OpenStudy (anonymous):

If the equation x^2-kx+3k=0 has exactly 1 solution for "x", which of the following could be "k"? 4 6 9 12 18

OpenStudy (anonymous):

b^2-4ac has to be 0

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

use the discriminant?

OpenStudy (anonymous):

solve it. you need one solution, so you need discriminant to be 0

OpenStudy (anonymous):

hmmmm..... but this equation is structured a little differently

OpenStudy (anonymous):

b^2-4ac=0 so (-k)^2-4*1*3k=0 k^2-12k=0

OpenStudy (anonymous):

k(k-12)=0 k=0 or k=12

OpenStudy (anonymous):

so k=12

OpenStudy (anonymous):

thank u. i'll need 2 remember this for future reference. thanx, bro

OpenStudy (anonymous):

generally... remember that

OpenStudy (anonymous):

thanx, anyway, bro

OpenStudy (anonymous):

 no matter k's and stuff